OpenStudy (anonymous):

Write the quadratic function in standard form. y= -(x+2)^2-7

OpenStudy (whpalmer4):

standard form is y = ax^2 + bx + c so, you need to expand that (x+2)^2 bit and arrange the results accordingly

OpenStudy (anonymous):

What do you do with the exponent 2?

OpenStudy (whpalmer4):

(x+2)(x+2) = x*2 + 2*x + 2*x + 2*2 = x^2 + 4x + 4 so y = -(x+2)^2 - 7 = - (x^2 + 4x + 4) - 7 = -x^2 - 4x - 4 - 7 = -x^2 - 4x - 11 @eliassaab got to watch for those leading - signs :-)

OpenStudy (anonymous):

\[(x+2)^2= x^2 + 4 x +4\\ y=-(x+2)^2-7= -x^2 -4 x -4-7= x^2 + 4 x -3\\ y= -x^2- 4 x -11 \]

OpenStudy (anonymous):

@whpalmer4, thanks for you remark. I am getting old to see the minus sign.